a^3-b^3=56 and ab=8 then what is the value of a-b. ?

Grade:10

1 Answers

Arun
6 years ago
Dear Amar
 
We have to find (a-b)
Now
(a-b)³ = a³ -b³ -3ab(a-b)
(a-b)³ = 56 -24 (a-b)
It is like 
x³ +24x -56 = 0
x = 2 it is true
Hence
a -b = 2
 
Regards
Arun (askIITians forum expert)